Mole fraction of solute in its 1molal solution is ________?

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
molality means the no. of moles of the solute in 1 kg of solvent (water). Here, molality=1, so there is 1 mole of solute in 1 kg or 1000 g of water.
 
as weight of water = 1000g and molar mass of water(H2O)=18g,
moles of water=1000/18=55.55
and we saw that moles of solute=1
so mole fraction =1/(1+55.55)=0.018.
 
Thanks,
Soumya Ranjan Mohanty.